OpenStudy (anonymous):

Solve. 1/x - 1/x+4 = 1/3

OpenStudy (anonymous):

solve for x? :)

OpenStudy (anonymous):

solve for x? :)

OpenStudy (anonymous):

it just says solve

OpenStudy (anonymous):

1/x- 1/x+4 -1/3 =0 now get the lcd which is 3x(x+4) youll get 3x+12 - 3x - x^2 + 4x =0 then -x^2 -3x + 4 = 0 x^2 +3x - 4 = 0 factoring (x+4)(X-1) =0 YOULL GET X = -4 AND X = 1

OpenStudy (anonymous):

get the same denominators and you'll get :\[= (3x^2 + 12x -3x -3x^2 -12x)/(3x)(x+4)\]\[= -3x/3x(x+4)\]\[= -1/(x+4)\] I guess if that's what the question wants :)